# Quasistatic approximation in neuromodulation

### v4

@@ 2.6 Adaptation of QSA for relaxing individual assumptions
Figure 4: Adaptations of QSA framework. A. A Fourier pipeline solves the E-field in response to an arbitrary stimulation waveform in dispersive tissue by decomposing the waveform into its frequency components, with the number of the components $n$ sufficient to capture the waveform spectrum. Each component is solved individually under QSA using frequency-specific tissue parameters. The pipeline can also be used for capacitive tissue and/or for obtaining full-wave solutions, for which individual field solutions are no longer under QSA. Symbols with hats indicate frequency domain variables. B. Iterative pipelines solve nonlinear systems. i. For electrical stimulation, tissue conductivities can be a nonlinear function of the E-field, and electrode current and voltage can be related by a non-linear electrode–tissue interface via the overpotential $V_{\mathrm{OP}}$ of the interface. ii. For magnetic stimulation using coils with nonlinear magnetic materials, the permeability is a function of the magnetic field intensity $\boldsymbol{H}$, and the primary E-field depends nonlinearly on the coil current. The gray boxes indicate that the computation is performed for a discrete time step. In some cases, information needs to be transferred from one step to the next, e.g., for capacitive electrode interface and hysteretic magnetic core.
@@ 3.2 Early and general development of QSA
Historically, the use of QSA was driven in part by the reliance on analytical or limited numerical solution methods, coupled with simplified geometries such as point electrodes or planar or spherical tissue structures [135, 136, 137, 39]. The work by Rush and Driscoll [135] is generally noteworthy for its early attempt to model and experimentally verify current flow models of transcranial electrical stimulation and serves as an example of the implicit adoption of QSA in neuromodulation. With the introduction of more advanced numerical techniques like FEM and increased computational resources, more intricate electrode and tissue geometries are implemented [138, 139, 140]. While a small minority of studies have applied non-QSA approaches, QSA remains almost universal at the stage of calculating tissue E-fields given its overall high accuracy and flexibility for adapting to non-QSA situations. Importantly, QSA makes targeting and dosimetry using E-field simulations conceptually simpler and more interpretable. Although improvement in computational power has allowed novel methods to handle some non-QSA simulations effectively, they may not provide significant improvement given that the accuracy of the field solutions are limited by the uncertainty of tissue conductivity.
Significant effort went into developing models of neuronal membrane polarization, action potential generation, and network stimulation (i.e., starting from predicted E-field through prediction of physiological outcomes), including analysis of the importance of neuronal morphology and biophysics [136, 141]. The four QSA assumptions are in the implicit background of these studies, when either calculating (e.g., FEM or analytical) or assuming (e.g., quasi-uniform) an imposed E-field. In some cases, these reports cite Plonsey and note “purely resistive medium” [142, 143, 39]. Thus, canonical studies establishing neuronal modeling methods implicitly propagated QSA (in a general sense) in neuromodulation by coupling the spatial and temporal components of the E-field separately to compartmental neuronal models. In contrast, non-QSA field modeling approaches would require adjustment for integration into most modeling pipelines, with the neural simulation modified to couple different temporal waveforms of the stimulation E-field to the neuronal compartments according to their location. As significant spatial variations in temporal waveforms are unlikely to be present on microscopic scales, non-QSA modeling approaches may only be necessary for very long fibers or large cells and so far were mostly used in studies evaluating the accuracy of QSA [50, 51, 3, 6, 5].
@@
3.3 Spinal cord stimulation, DBS, and point source (microelectrode) stimulation
@@
Models of tDCS adopted the four QSA assumptions. Such efforts started with concentric sphere models [160, 161] and were extended to smooth anatomy rendered from magnetic resonance imaging (MRI) using computer aided design [162] and contemporary MRI-derived “gyri-precise” anatomy [139]. Under direct current, assumptions A3 and A4 seem evident since the stimulation is not changing with time [163]. In tDCS studies, QSA may not even be mentioned but is implied by the reduced governing equations (see section 2.3.2). Nonetheless, open questions on modeling methods remain, such as which tissue conductivities are appropriate (given that conductivities are rarely measured using direct currents) and how they may change with current or time as suggested by the reduction in overall resistance [159]. There have been some validation studies of predicted tDCS current flow [164, 165, 166], although more extensive evidence may come from validation of tACS, which is considered to be computationally analogous.
@@
A recent study directly accessed the QSA errors for tACS, with specific clarification of each assumption [6]. QSA-induced E-field errors were analyzed as a function of frequency in the 10 Hz to 100 MHz range, necessitating an implicit relaxation of assumption A4. Subsequently, assumption A3 was relaxed to examine the impact of tissue capacitance on the error. The reference solution employed a general case of the full set of Maxwell’s equations (i.e., assumption A1 was also omitted, and wave effects were considered). When comparing the results involving tissue capacitance (without assumption A3) to this general solution, the limitation of assumption A1 was identified in the MHz range. However, assumption A3 was observed to introduce up to a 20% error in E-field the brain at 10 Hz.
@@ 4 Conclusions and recommendations for communication of QSA in neuromodulation
    Parameters such as assumed tissue conductivity values, boundary conditions (see sections 2.3 and 2.4), and other model and solver settings should be considered and reported for rigor and reproducibility of the models. The errors due to uncertainties in the parameter values or deviations from any of the assumptions A1–A4 should be estimated if these errors may influence the validity of the conclusions of the study in question.

### v5

@@ 2.6 Adaptation of QSA for relaxing individual assumptions
Figure 4: Adaptations of QSA framework. A. A Fourier pipeline solves the E-field in response to an arbitrary stimulation waveform in dispersive tissue by decomposing the waveform into its frequency components, with the number of the components $n$ sufficient to capture the waveform spectrum. Each component is solved individually under QSA using frequency-specific tissue parameters. The pipeline can also be used for capacitive tissue and/or for obtaining full-wave solutions, for which individual field solutions are no longer under QSA. Symbols with hats indicate frequency domain variables. B. Iterative pipelines solve nonlinear systems. i. For electrical stimulation, tissue conductivities can be a nonlinear function of the E-field, and electrode current and voltage can be related by a non-linear electrode-tissue interface via the overpotential $V_{\mathrm{OP}}$ of the interface. ii. For magnetic stimulation using coils with nonlinear magnetic materials, the permeability is a function of the magnetic field intensity $\boldsymbol{H}$, and the primary E-field depends nonlinearly on the coil current. The gray boxes indicate that the computation is performed for a discrete time step. In some cases, information needs to be transferred from one step to the next, e.g., for capacitive electrode interface and hysteretic magnetic core.
@@ 3.2 Early and general development of QSA
Historically, the use of QSA was driven in part by the reliance on analytical or limited numerical solution methods, coupled with simplified geometries such as point electrodes or planar or spherical tissue structures [120, 121, 122, 36]. The work by Rush and Driscoll [120] is generally noteworthy for its early attempt to model and experimentally verify current flow models of transcranial electrical stimulation and serves as an example of the implicit adoption of QSA in neuromodulation. With the introduction of more advanced numerical techniques like FEM and increased computational resources, more intricate electrode and tissue geometries are implemented [123, 124, 125]. While a small minority of studies have applied non-QSA approaches, QSA remains almost universal at the stage of calculating tissue E-fields.
Significant effort went into developing models of neuronal membrane polarization, action potential generation, and network stimulation (i.e., starting from predicted E-field through prediction of physiological outcomes), including analysis of the importance of neuronal morphology and biophysics [121, 126]. The four QSA assumptions are in the implicit background of these studies, when either calculating (e.g., FEM or analytical) or assuming (e.g., quasi-uniform) an imposed E-field. In some cases, these reports cite Plonsey and note “purely resistive medium” [127, 128, 36]. Thus, canonical studies establishing neuronal modeling methods implicitly propagated QSA (in a general sense) in neuromodulation.
@@
3.3 Spinal cord stimulation, DBS and point source (microelectrode) stimulation
@@
Models of tDCS adopted the four QSA assumptions. Such efforts started with concentric sphere models [145, 146] and were extended to smooth anatomy rendered from magnetic resonance imaging (MRI) using computer aided design [147] and contemporary MRI-derived “gyri-precise” anatomy [124]. Under direct current, assumptions A3 and A4 seem evident since the stimulation is not changing with time [148]. In tDCS studies, QSA may not even be mentioned but is implied by the reduced governing equations (see section 2.3.2). Nonetheless, open questions on modeling methods remain, such as which tissue conductivities are appropriate (given that measurements of conductivities rarely use dc test currents) and how they may change with current or time as suggested by the reduction in overall resistance [144]. There have been some validation studies of predicted tDCS current flow [149, 150, 151], although more extensive evidence may come from validation of tACS, which is considered to be computationally analogous.
@@
A recent study directly accessed the QSA errors for tACS, with specific clarification of each assumption [6]. QSA-induced E-field errors were analyzed as a function of frequency in the 10 Hz to 100 MHz range, necessitating an implicit relaxation of assumption A4. Subsequently, assumption A3 was relaxed to examine the impact of tissue capacitance on the error. The reference solution employed a general case of the full set of Maxwell’s equations (i.e., assumption A1 was also omitted, and wave effects were considered). When comparing the results involving tissue capacitance (without assumption A3) to this general solution, the limitation of assumption A1 was identified in the MHz range. However, assumption A3 was observed to introduce up to a 20% error in the brain at 10 Hz.
@@ 4 Conclusions and recommendations for communication of QSA in neuromodulation
    Parameters such as assumed tissue conductivity values, boundary conditions (see sections 2.3.2 and 2.3.3), and other model and solver settings should be considered and reported for rigor and reproducibility of the models. The errors due to uncertainties in the parameter values or deviations from any of the assumptions A1–A4 should be estimated if these errors may influence the validity of the conclusions of the study in question.